A highway construction company is under contract to build a new roadway through a scenic area and two rural towns in Colorado. The road is expected to cost $10,000,000, with annual upkeep estimated at $150,000 per year. Additional income from tourist (Benefits) of $900,000 per year is estimated. The road is expected to have a useful commercial life of 15 years. The discount rate to evaluate this project is 6% per year.

Hint: Disbenefits = $0 and Annual Upkeep is equivalent to Annual M&O

Based on the data provided the Conventional B/C is equal to:

(All the alternatives presented below were calculated using compound interest factor tables including all decimal places)

  1. Conventional B/C = 0.523
  1. Conventional B/C = 1.311
  1. Conventional B/C = 0.763
  1. Conventional B/C = 1.023
Answer #1

Conventional B/C = (AW of benefits - AW of disbenefits) / (AW of initital cost + AW of O&M - AW of salvage cost)

Conventional B/C = 900000 / (10000000 * (A/P,6%,15) + 150000)

= 900000 / (10000000 * 0.1030 + 150000)

= 900000 / 1180000

= 0.762712 ~ 0.763

Third option is correct answer